OpenStudy (anonymous):

Please help with this vectors question Write a vector equation and a scalar equation of the plane. Parallel to the plane 3x - 9y + z - 12 = 0 and including the point (-3, 7, 1) Answer at the back of the textbook: [x, y, z] = [-3, 7, 1] + s[3, 1, 0] + t[1, 0, -3] 3x - 9y + z + 71 = 0 Please show all work! Thanks :)

OpenStudy (anonymous):

Is this correct? Starts from the top.

OpenStudy (kinggeorge):

It looks correct to me.

OpenStudy (anonymous):

But is there anything that I can change so its not "copying"?

OpenStudy (anonymous):

and isnt there supposed to be 3 terms instead of 2 for the final answer?

OpenStudy (anonymous):

like [x,y,z]=[a,b,c]+s[d,e,f]+t[g,h,i] instead of, [x,y,z]=[a,b,c]+s[d,e,f]

OpenStudy (anonymous):

@KingGeorge ?

OpenStudy (anonymous):

I did up to the scalar equation which is, 3x - 9y + z + 71 = 0 I just need help with the vector equation now

OpenStudy (kinggeorge):

OK, there are supposed to be 3 terms instead of two. Whoever wrote on that paper gave you the first two terms. Now you just need the 3rd. To do that, find another point in the plane.

OpenStudy (anonymous):

Ok, so can it be [-1, 2, 5]?

OpenStudy (anonymous):

By the way, what point did they use?

OpenStudy (kinggeorge):

You need to make sure that it works with the equation 3x-9y+z+71=0, so (-1, 2, 5) would not work. To make this work, choose an x value, and a y value, and solve for z. They used the point (2, 1, -68)

OpenStudy (anonymous):

How about, [3, -10, 10] ?

OpenStudy (anonymous):

**[3, 10, 10]

OpenStudy (kinggeorge):

Let x=3, y=10. That means you have \[3(3)-9(10)+z+71=-81+71+z=-10+z\]So that point should work.

OpenStudy (kinggeorge):

Now, what's the vector from (-3, 7, 1) to that new point?

OpenStudy (anonymous):

[3, 10, 10] - [-3, 7, 1] [-3-3, 7-10, 1-10] [-6, -3, -9] ?

OpenStudy (kinggeorge):

Just about. It should be [6, -3, -9]. This is the second vector you need for your equation.

OpenStudy (anonymous):

-3-3=-6

OpenStudy (kinggeorge):

It should be \(3-(-3)=6\).

OpenStudy (anonymous):

why though? I thought it was right - left?

OpenStudy (kinggeorge):

It's the second point - the first point. We're using the second point as (3, 10, 10) and the first as (-3, 7, 1). You even typed the first part correctly. Now that I think about it, all the negatives should be reversed. So your vector should be (6, 3, 9)

OpenStudy (anonymous):

And what about the previous question, The points were [-1, 0, 0] and [-1, 3, 4] and the point that was given was [-1, -2, -5]

OpenStudy (kinggeorge):

The point was actually [-1, -2, 5], and the vector (for the second one) was [-1,3,4]-[-1,-2,5]=[0,3-(-2),4-5]=[0,5,-1]

OpenStudy (anonymous):

Oh yea my fault

OpenStudy (anonymous):

Ok, so the answer is, [x, y, z] = [-3, 7, 1] + ?

OpenStudy (kinggeorge):

The first vector that was written on the paper, which is [5, -6, 69] and the vector we just derived here which is [6, 3, 9].

OpenStudy (anonymous):

which is multiplied by s?

OpenStudy (kinggeorge):

It doesn't matter. As long as one is multiplied by s, and the other by t, you should be fine.

OpenStudy (anonymous):

I see, THANKS AGAIN! Sorry for bothering you so much :P Have a good day! :)